### v1

# Phase transitions and scale invariance in topological Anderson insulators

Abstract
We investigate disordered-driven transitions between trivial and topological insulator (TI) phases in two-dimensional (2D) systems. Our study primarily focuses on the BHZ model with Anderson disorder, while other standard 2DTI models exhibit equivalent features. The analysis is based on the local Chern marker (LCM), a local quantity that allows for the characterization of topological transitions in finite and disordered systems. Our simulations indicate that disorder-driven trivial to topological insulator transitions are nicely characterized by $\mathcal{C}_{0}$, the disorder averaged LCM near the central cell of the system. We show that $\mathcal{C}_{0}$ is characterized by a single-parameter scaling, namely, $\mathcal{C}_{0}(M,W,L)\equiv\mathcal{C}_{0}(z)$ with $z=[W^{\mu}-W_{c}^{\mu}(M)]L$, where $M$ is the Dirac mass, $W$ is the disorder strength and $L$ is the system size, while $W_{c}(M)\propto\sqrt{M}$ and $\mu\approx 2$ stand for the critical disorder strength and critical exponent, respectively. Our numerical results are in agreement with a theoretical prediction based on a first-order Born approximation (1BA) analysis. These observations lead us to speculate that the universal scaling function we have found is rather general for amorphous and disorder-driven topological phase transitions.
Introduction.– Topological insulators are a fascinating new class of materials characterized by unique electronic properties [1, 2, 3]. Notable among their attributes is the appearance of symmetry-protected spin-polarized edge (or surface) states that contrast with the insulating bulk region [4, 5]. These materials find diverse applications in spintronics, encompassing spin-to-charge conversion (SCC) devices [6], memory read-out, field-effect transistors in quantum computing [7], and demonstrate promising potential in preparing electrodes for photodetectors [8] and other applications [9, 10, 11].
Since the unique features of such materials manifest only in their topological phase, there is great interest in a deeper understanding of their robustness. The theory predicts that smooth deformations of the system band structure do not change its topological properties, provided the topological gap does not close [1, 2]. In pristine systems, topological invariants [4, 12] have been used to characterize the topological phase and to study the trivial to topological transitions caused by magnetic fields [13], electric fields [14], mechanical strain [15], temperature variations [16], doping, and interactions. The understanding of disordered systems is less clear since the standard topological invariants rely on translation symmetry.
A systematic study of topological phase transitions [17] has put forward a scaling procedure for inversion-symmetric topological insulators, successfully identifying critical and fixed points under scaling of the curvature function. Similar results were obtained for other topological systems, such as in static and periodically driven Kitaev chains [18] and in higher-order band crossings [19]. Topological phase transitions have also been studied using the LCM as a function of a mass parameter in the Haldane model [20].
Of particular interest are disorder-driven, such as the Anderson Topological Insulators (TAI) [21, 22], and amorphous [23, 24, 25, 26] topological systems, where one has to rely on local topological invariants, such as the Local Charn Marker (LCM) [27] and others [28, 29, 30]. Various studies have identified TAI systems, including disordered Haldane models and Kane-Mele systems [31, 32], bilayer crystals [33], three-dimensional systems [34], and others [35, 36]. Experimental realizations include heterostructures [37], disordered atomic wires [38], among others [39, 40]. However, a study of scaling and critical exponents in disorder-driven topological phase transitions is still lacking.
Figure 1: (a,b) Band structure of the BHZ Hamiltonian for a pristine ribbon of width $L_{y}=600$ nm with (a) $M=2$ meV, and (b) $M=-10$ meV. (c) Conductance $G$ averaged over $N=100$ disorder realizations, as a function of the chemical potential $\varepsilon_{F}$ and disorder strength $W$. The yellow region highlights the $G=G_{0}$ quantized plateau corresponding to the TAI phase. (d-f) Current density (black lines and arrows) and spin polarization density (color coded with spin up in red, and spin down in blue) for a single disordered realization with $L_{x}=L_{y}=100$ nm and (d) $W=180$ meV, (e) $W=400$ meV, and (f) $W=897$ meV.
In this Letter, we introduce the disorder averaged local Chern marker at the center of the system $\mathcal{C}_{0}$ to characterize trivial to topological phase transitions driven by the disorder. For finite-size systems, $\mathcal{C}_{0}$ varies smoothly as a function of the disorder strength $W$, as defined by the Anderson disorder model, allowing for a scale invariance investigation with the system size $L$. Here, we consider the BHZ model [41], a standard model used to study 2DTIs, while equivalent results for the Haldane [42] and Kane-Mele [43] models are shown in the Supplemental Materials [44]. Our simulations indicate that $\mathcal{C}_{0}$ scales as $\mathcal{C}_{0}(W,L)\equiv\mathcal{C}_{0}(z)$ with $z=(W^{\mu}-W_{c}^{\mu})L$, where $W_{c}$ is the critical disorder strength and $\mu\approx 2$ is the critical exponent in the trivial to TAI phase transition. These results are further supported by a theory based on the scaling analysis of first-order Born approximation (1BA) results. These observations lead us to speculate that our findings are universal for disorder-driven topological phase transition.
Model and Methods.– The BHZ Hamiltonian [41] is one of the standard models to describe 2D TIs in the AII symmetry class of the Altland-Zirnbauer 10-fold way classification [45, 46]. Considering only the majority spin up block, the BHZ Hamiltonian reads
$H=C-Dk^{2}+A\bm{k}\cdot\bm{\sigma}+(M-Bk^{2})\sigma_{z},$ (1)
[…]
Figures 1(a) and 1(b) show the band structure of the BHZ Hamiltonian in a ribbon geometry of width $L_{y}=600$ nm and periodic along $x$ for $M=+2$ meV and $M=-10$ meV, respectively. In both cases, we consider pristine systems ($W=0$) to emphasize that in the trivial insulator case, there is a band gap of $\approx 2M=4$ meV, while in the nontrivial regime ($M<0$) the gap is filled with topologically protected helical edge states, that are responsible for a quantized conductance of $G_{0}=e^{2}/h$. As shown in Ref. [22], and revised in the Supplemental Materials [44], within the first-order Born approximation (1BA) the Anderson disorder ($W\neq 0$) renormalizes the mass term as
$M\rightarrow\widetilde{M}\approx M-\alpha(M)W^{2}.$ (2)
This important result reveals that disorder can drive the system from a trivial regime ($M=2$ meV, $W=0$) into a Topological Anderson Insulator (TAI) regime with an effective mass $\widetilde{M}<0$. Figure 1(c) shows the conductance $G$ for the disordered BHZ model as a function of chemical potential $\varepsilon_{F}$ and disorder intensity $W$. The large parameter space region of the quantized conductance $G=G_{0}$ (in yellow) characterizes a TAI phase. The panels Fig. 1(d-e) illustrate the current and spin density for a smaller system ($L=100$ nm) for different disorder intensities, and chemical potential $\varepsilon_{F}=0$. For small $W$, as shown in Fig. 1(d), the system is already near a TAI phase and the current and spin densities show a tendency to form helical edge states. Fig. 1(e) shows that as the disorder strength $W$ is increased, the system reaches the TAI phase and the helical edge states become well-defined. For even larger $W$ values, comparable with the bandwidth, the topological protection is destroyed and the states become localized, as shown by Fig. 1(f), indicating the onset of Anderson localization.
Local Chern Marker.– As discussed in Ref. [21], in the trivial regime the conductance is dominated by bulk transport, and the scaled conductance $\sigma=GL_{x}/L_{y}$ decays with $W$ as a universal function, independent of the ribbon width $L_{y}$. In contrast, at the TAI phase, $G=G_{0}$ forms a quantized plateau, see Fig. 1(c). Therefore, instead of using $G$, we find it convenient to use local markers to investigate the normal-TAI transition, its scale invariance properties, and critical exponents. As shown in Ref [49], the local Chern marker can be written as [50]
[…]
where the integral is taken over a unit cell of area $A_{c}$ centered at $\bm{R}$, the sum runs over the states $\beta$ that define the basis for the model Hamiltonian (e.g, $E_{1}/H_{1}$ orbitals for the BHZ model), $\hat{x}$ and $\hat{y}$ are position operators, and $\hat{P}$ is the projector over occupied states, defined as
$\hat{P}=\sum_{\ell}^{\epsilon_{\ell}<E_{F}}\outerproduct{\psi_{\ell}}{\psi_{\ell}}.$ | . (4)
Here |$\ket{\psi_{\ell}}$ and $\epsilon_{\ell}$ are the eigenenergies and eigenstates of the Hamiltonian $H$ for a finite, nonperiodic system. It has been shown that $\mathcal{C}(\bm{R})$ recovers the quantized bulk Chern number $C$ as one restores the lattice periodicity for clean systems [30], while for finite systems $\mathcal{C}(\bm{R})\approx C$ for $\bm{R}\approx 0$, near the center of the sample [49, 20].
In the presence of Anderson disorder, the local marker $\mathcal{C}(\bm{R})$ acquires fluctuations. Consequently, the analysis of the LCM requires a statistical approach. Here, we focus on the average $\mathcal{C}(\bm{R})$, that can be taken over disorder realizations and/or spatial averages. For large systems, this can be an intensive numerical task. Therefore, we have implemented two approaches. First, to calculate the full $\mathcal{C}(\bm{R})$ map as a function of $\bm{R}$ for a single disorder sample, we use a brute force approach, where the finite size system is fully diagonalized, $H\ket{\psi_{\ell}}=\epsilon_{\ell}\ket{\psi_{\ell}}$, such that the operators $\hat{P}$, $\hat{x}$, $\hat{y}$ can be explicitly calculated to yield $\mathcal{C}(\bm{R})$ from Eq. (3). In the second approach, we define the disorder average local marker $\mathcal{C}_{0}$ as the average of $\mathcal{C}(\bm{R}=0)$ over many disorder realizations, namely, $\mathcal{C}_{0}\equiv\langle\mathcal{C}(\bm{R}=0)\rangle$. We compute $\mathcal{C}_{0}$ employing the efficient implementation developed in Ref. [51] based on the Kernel Polynomial Method (KPM). The KPM allows us to estimate $\hat{P}\ket{\varphi}$ for an arbitrary state |$\ket{\varphi}$ in terms of an efficient expansion of Chebyshev polynomials.
Figure 2 shows $\mathcal{C}(\bm{R})$ as a function of $\bm{R}=(X,Y)$ for a single disorder realization for increasing disorder strengths $W$. Panel (i) corresponds to a small $W$ for which the system is in the trivial regime, that is characterized by a small LCM near the sample center, that is, $\mathcal{C}({\bm{R}}\approx 0)\approx 0$. Panel (ii) considers a moderate disorder intensity $W$, yielding $\mathcal{C}({\bm{R}}\approx 0)\approx 0.7$, thus approaching a TAI regime, which is fully achieved for larger $W$, as shown in panel (iii), where $\mathcal{C}({\bm{R}}\approx 0)\approx 1.0$. Notice that panels (ii) and (iii) of Fig. 2 correspond to the same values of $W$ of Figs. 1(d) and 1(e). Interestingly, up to Fig. 2(iii) or Fig. 1(e), we find that the Anderson disorder has driven the system into a TAI phase, but there are no clear signs of Anderson localization, since the maps are overall smooth, apart from small fluctuations that cause some blurriness. However, for even larger $W$, see Fig. 2(iv) or Fig. 1(f), Anderson localization emerges as strong fluctuations over the full map $\mathcal{C}(\bm{R})$. In all cases, we see that $\mathcal{C}(\bm{R})$ takes negative values at the edges, which fulfills the constraint $\sum_{\bm{R}}\mathcal{C}(\bm{R})=0$ [49]. Indeed, this is an undesirable consequence of the approximations that take place in the derivation of the local marker [50], which, for now, we leave as an issue to be discussed in future works.
Figure 2: Local Chern marker $\mathcal{C}(\bm{R})$ for a representative single-disorder realization as a function of $\bm{R}=(X,Y)$ for a system of size $L_{x}=L_{y}=100$ nm, $M=+2$ meV, and disorder strengths (i) $W=25$ meV, (ii) $W=180$ meV, (iii) $W=400$ meV and (iv) $W=897$ meV. The panels (ii) to (iv) correspond to the current and spin densities shown in Fig. 1(e-f).
Phase transition and scaling.– Figure 3 shows the disorder averaged local Chern marker $\mathcal{C}_{0}$ as a function of $W$ for different system sizes in a square geometry $L_{x}=L_{y}\equiv L$ and a fixed $M=+2$ meV. As $W$ increases, all cases first display a transition from a trivial regime with $\mathcal{C}_{0}\approx 0.1$ to a nontrivial regime characterized by $\mathcal{C}_{0}\approx 1$. For very large values of $W$, exceeding the band width ($E_{\rm max}\sim|B|k_{\rm max}^{2}=8|B|/a^{2}\approx 610$ meV for $a=3$ nm) of the pristine BHZ Hamiltonian, $\mathcal{C}_{0}$ collapses back to $\mathcal{C}_{0}\rightarrow 0$. This is the transition seen from panels (iii) to (iv) in Fig. 2, which is due to the emergence of Anderson localization. Such behavior is not surprising, since in such a strong disorder regime, it hardly makes sense to consider the system as topological. Figure 3(a) clearly shows that the second transition, $\mathcal{C}_{0}=1\rightarrow 0$, does not dependent on $L$, whereas the first one, $\mathcal{C}_{0}\approx 0.1\rightarrow 1$, does. This dependence is what we analyze next.
We obtain a single-parameter universal scaling function for the trivial-TAI phase transition by expressing ${\cal C}_{0}$ in terms of the parameter $z$, that combines the disorder strength $W$ and the system size $L$ according to
$\mathcal{C}_{0}(W,L)\rightarrow\mathcal{C}_{0}(z),$ (5)
with
$z=\dfrac{W^{\mu}-W_{c}^{\mu}}{W_{\rm max}^{\mu}}\cdot\dfrac{L}{L_{\rm max}},$ (6)
as shown in Fig. 3(b). Here, $W_{c}$ is the critical point where the curves corresponding to systems with different sizes $L$ cross, and $\mu$ is the critical exponent. To make $z$ dimensionless, we introduce $L_{\rm max}=250$ nm, the largest simulated system size, and $W_{\rm max}^{\mu}=1200$ meV, the largest considered disorder strength, in the denominators of $z$. We find numerically that the optimal scaling in Fig. 3(b) occurs for $\mu\approx 2$ and $W_{c}\approx 86$ meV [dashed line in Fig. 3(a)]. Next, to understand these values for $\mu$ and $W_{c}$, we analyze the phase diagram and scaling properties of the $\mathcal{C}_{0}$ as a function of both the mass $M$ and disorder intensity $W$.
Figure 3: (a) Disorder averaged local Chern marker (400 disorder realizations), as a function of $W$ for different $L_{x}=L_{y}=L$ showing a trivial-TAI transition, $\mathcal{C}_{0}\rightarrow 0$ to 1, followed by the trivial Anderson localization as $W$ becomes larger than the band width. (b) Optimal scaling of the maker showing that all curves in (a) fall into a universal trend for the trivial-TAI transition region upon the scaling $W\rightarrow z=(W^{\mu}-W_{c}^{\mu})L/W_{\rm max}^{\mu}L_{\rm max}$, with $\mu\approx 2$, and $W_{c}\approx 86$ meV. The points indicated by the arrows (i) to (iv) correspond to the ${\cal C}({\bm{R}})$ color maps shown in Fig. 2.
Phase diagram.– The local Chern marker can be used to define a topological phase diagram. For the BHZ model, we now consider $\mathcal{C}_{0}$ in terms of the mass $M$ and disorder strength $W$, as shown in Fig. 4, for two system sizes. In both cases the color code labels the value of the local marker $\mathcal{C}_{0}$ calculated for 400 disorder realizations. The phase diagrams on panels (a) and (b), show the following phases: trivial, TI, TAI, and Anderson localization (AL). As discussed above, the AL phase occurs when the disorder strength exceeds the band width $W>E_{\rm max}\approx 610$ meV.
We recall that, for $W=0$, the bulk Chern number indicates that a TI phase occurs for $M<0$, whereas Figs. 4(a) and 4(b) show that the trivial to TI transition, characterized by the local marker $\mathcal{C}_{0}$, depends upon the finite system size $L$. In a pristine 2DTI, the topological edge states are gapless and do not hybridize, provided the system size $L$ is much larger than the edge states penetration depth $\xi=A/|M|$, i.e., for $A/|M|\ll L$ (with $M<0$ and $W=0$). We extend this gapless condition to the disordered case by making $M\rightarrow\widetilde{M}$ to obtain an expression for the phase boundary using $\widetilde{M}$ from 1BA in Eq. (2). Hence, from the gapless condition $A/|\widetilde{M}|\ll L$ and assuming $\widetilde{M}\leq 0$, we obtain
$W_{\rm gapless}(M)=\sqrt{\dfrac{M+A/L}{\alpha(M)}}\approx\sqrt{\dfrac{M+A/L}{\alpha_{0}}}.$ (7)
[…]
Figure 4: (a-b) Phase diagrams calculated using the local Chern marker $\mathcal{C}_{0}$ as a function of the BHZ mass $M$ and disorder strength $W$ for (a) $L=50$ nm and (b) 100 nm. The dashed white line marks the phase boundary between the trivial and topological phases (TI and TAI) obtained from the gapless condition $A/|\widetilde{M}|\ll L$. The dot-dashed line marks the boundary where $W$ matches the band width of the clean system, above which the Anderson localization phase takes place. (c) All data from panels (a) and (b), and additional data for a system of $L=150$ nm, fall into a universal curve for the trivial to TAI phase transition upon the scaling $(M,W,L)\rightarrow z=[W^{\mu}-W_{c}^{\mu}(M)]L/W^{\mu}_{\rm max}L_{\rm max}$.
The data shown in Figs. 4(a) and (b) for all $M$ and $W$, and additional data for $L=150$ nm (not shown), can be combined into a single universal function that characterizes the trivial to TAI phase transition shown in Fig. 4(c). The scaling used in this case is a direct extension of the one presented in Eqs. (5) and (6), where now $W_{c}\rightarrow W_{c}(M)$ depends upon $M$. To obtain an expression for $W_{c}(M)$ and to understand the value $\mu\approx 2$ of the critical parameters, let us first consider the clean regime ($W=0$), to find the scaling law for the trivial to TI phase transition as a function of $M$. Here, we consider an argument similar to the one used in Ref. [20] for the Haldane model, as follows. In the topological regime, the only relevant length scales are $\xi=A/|M|$ and $L$, as introduced above. Therefore, one can expect that the topological phase transition scaling with the system size $L$ should be a universal function of $L/\xi\propto ML$ (assuming $A$ is constant). Thus, we justify that
$\mathcal{C}_{0}(M,L)\equiv\mathcal{C}_{0}(ML),$ (8)
Next, to extend this scaling to finite $W$, we replace $M\rightarrow\widetilde{M}$ using the 1BA expression from Eq. (2) with $\alpha(M)\approx\alpha_{0}$, such that the universal function argument now reads
$\widetilde{M}L=(M-W^{2})L=-\alpha_{0}(W^{2}-W_{c}^{2})L,$ (9)
where $W_{c}^{2}=M/\alpha_{0}$ is the critical point and marks the phase boundary in the thermodynamic limit ($L/\xi\rightarrow\infty$), where $\widetilde{M}=0$ and $W_{\rm gapless}=W_{c}$. Indeed, for $M=2$ meV, this expression yields $W_{c}\approx 86$ meV, thus matching the value found in Fig. 3. Moreover, apart from constant factors, the expression for $\widetilde{M}L$ in Eq. (9) matches $z$ given by Eq. (6) with $\mu=2$, in agreement with the previously numerically obtained value.
Conclusions.– We have investigated the scale invariance of the local Chern marker $\mathcal{C}_{0}$ applied to topological Anderson insulators (TAI). The smooth profile of $\mathcal{C}_{0}$ throughout the phase transition allows us to characterize the scale invariance, obtain critical parameters, and universal scaling functions. Our analysis uses the disordered BHZ model as a canonical example of TI and TAI phases. The simplicity of the model allows for an analytical justification of the obtained numerical results. We like to stress that the disordered Haldane [42] and Kane-Mele [43] models exhibit an identical scaling and universal behavior, as we have shown in the Supplemental Material [44]. This observation supports the statement that our findings are quite generic. Indeed, as long as the first-order Born approximation remains sufficiently accurate, we expect that the universal function scales with $\sim W^{2}L$, while deviations from this behavior should occur only beyond the validity range of the 1BA. We believe that our findings shed light on an approach to investigating scaling properties of topological phase transitions for disordered and amorphous systems.
Note added: shortly before we finished this Letter, Ref. [52] - that has a large overlap with our study - appeared in the arXiv.
Acknowledgments.– This work was supported by the Brazilian funding agencies CNPq, CAPES, FAPERJ, and FAPEMIG (grant PPM-00798-18). The authors acknowledge useful conversation with C. Beenakker, A. Akhmerov, and C. P. Orth.