### v2

# Phase transitions and scale invariance in topological Anderson insulators

Abstract
We investigate disorder-driven transitions between trivial and topological insulator (TI) phases in two-dimensional (2D) systems. Our study primarily focuses on the Bernevig-Hughes-Zhang (BHZ) model with Anderson disorder, while other standard 2DTI models exhibit equivalent features. The analysis is based on the local Chern marker (LCM), a local quantity that allows for the characterization of topological transitions in finite and disordered systems. Our simulations indicate that disorder-driven trivial to topological insulator transitions are nicely characterized by $\mathcal{C}_{0}$, the disorder-averaged LCM near the central cell of the system. We show that $\mathcal{C}_{0}$ is characterized by a single-parameter scaling, namely, $\mathcal{C}_{0}(M,W,L)\equiv\mathcal{C}_{0}(z)$, with $z=[W^{\mu}-W_{c}^{\mu}(M)]L$, where $M$ is the Dirac mass, $W$ is the disorder strength, and $L$ is the system size, while $W_{c}(M)\propto\sqrt{M}$ and $\mu\approx 2$ stand for the critical disorder strength and the critical exponent, respectively. Our numerical results are in agreement with a theoretical prediction based on a first-order Born approximation analysis. These observations lead us to speculate that the universal scaling function we have found is rather general for amorphous and disorder-driven topological phase transitions.
Introduction.– Topological insulators are a fascinating new class of materials characterized by unique electronic properties [1, 2, 3]. Notable among their attributes is the appearance of symmetry-protected spin-polarized edge (or surface) states that contrast with the insulating bulk region [4, 5]. These materials find diverse applications in spintronics, encompassing spin-to-charge conversion devices [6], memory read-out, and field-effect transistors in quantum computing [7], and they demonstrate promising potential in preparing electrodes for photodetectors [8] and other applications [9, 10, 11].
Since the unique features of such materials manifest only in their topological phase, there is great interest in a deeper understanding of their robustness. The theory predicts that smooth deformations of the system’s band structure do not change its topological properties, provided the topological gap does not close [1, 2]. In pristine systems, topological invariants [4, 12] have been used to characterize the topological phase and to study the trivial to topological transitions caused by magnetic fields [13], electric fields [14], mechanical strain [15], temperature variations [16], doping, and interactions [17]. The understanding of disordered systems is less clear since the standard topological invariants rely on translation symmetry.
A systematic study of topological phase transitions [18] has put forward a scaling procedure for inversion-symmetric topological insulators, successfully identifying critical and fixed points under scaling of the curvature function. Similar results have been obtained for other topological systems, such as in static and periodically driven Kitaev chains [19] and in higher-order band crossings [20]. Topological phase transitions have also been studied using the local Chern marker (LCM) as a function of a mass parameter in the Haldane model [21].
Figure 1: (a,b) Band structure of the BHZ Hamiltonian for a pristine ribbon of width $L_{y}=600$ nm with (a) $M=2$ meV, and (b) $M=-2$ meV. (c) Conductance $G$ averaged over $N=100$ disorder realizations, as a function of the Fermi energy $\varepsilon_{F}$ and the disorder strength $W$. The yellow region highlights the $G=G_{0}$ quantized plateau corresponding to the TAI phase. (d)-(f) Current density (black lines and arrows) and spin polarization density (color coded with spin up in red, and spin down in blue) for a single-disorder realization with $L_{x}=L_{y}=100$ nm and (d) $W=180$ meV, (e) $W=400$ meV, and (f) $W=897$ meV.
Of particular interest are disorder-driven, such as the topological Anderson insulators (TAI) [22, 23], and amorphous [24, 25, 26, 27] topological systems, where one has to rely on local topological invariants, such as the LCM [28, 29, 30]. Recent studies have extended the two-dimensional LCM put forward by Bianco-Resta [29] to odd dimensions [31], defined a local Chern marker for periodic systems [32, *StraWBerryPy] and finite temperatures [34], and established conditions for their robustness against disorder [35], opening new paths for future investigations. Various studies have identified TAI systems, including disordered HgTe quantum wells [36, 37, 38, 39], Haldane models and Kane-Mele systems [40, 41], bilayer crystals [42], three-dimensional systems [43], four-dimensional artificial lattices [44], higher-order TAI in Sierpiński lattices [45], half-filled Haldane models with extended Hubbard interactions [46], the Haldane model with binary disorder [47], and others [48, 49], where nontrivial phases arise in areas of the phase diagram where the clean limit is topologically trivial. The disordered topological phase can also be classified by investigating how the conductance approaches the quantization plateau [50, 51]. Experimental realizations include heterostructures [52] and disordered atomic wires [53], among others [54, 55]. However, a study of scaling and critical exponents in disorder-driven topological phase transitions is still lacking.
In this Letter, we introduce the disorder-averaged local Chern marker at the center of the system $\mathcal{C}_{0}$ to characterize trivial to topological phase transitions driven by the disorder. For finite-size systems, $\mathcal{C}_{0}$ varies smoothly as a function of the disorder strength $W$, as defined by the Anderson disorder model, allowing for a scale invariance investigation with the system size $L$. Our findings are based on the study of three seminal models for two-dimensional (2D) topological insulators (TIs), namely, the Bernevig-Hughes-Zhang (BHZ) [56], Haldane [57], and Kane-Mele [58] models. Here we focus on the BHZ model, while in order to avoid unnecessary repetitions, equivalent results for the Haldane and Kane-Mele models are shown in the Supplemental Materials [59]. In all cases, our simulations indicate that $\mathcal{C}_{0}$ scales as $\mathcal{C}_{0}(W,L)\equiv\mathcal{C}_{0}(z)$, with $z=(W^{\mu}-W_{c}^{\mu})L$, where $W_{c}$ is the critical disorder strength and $\mu\approx 2$ is the critical exponent in the trivial to TAI phase transition. These results are further supported by a theory based on the scaling analysis of first-order Born approximation (1BA) results. These observations lead us to speculate that our findings are universal for disorder-driven topological phase transitions.
Model and methods.– The BHZ Hamiltonian [56] is one of the standard models to describe 2D TIs in the AII symmetry class of the Altland-Zirnbauer tenfold way classification [60, 61]. Considering only the majority spin-up block, the BHZ Hamiltonian reads as
$H=C-Dk^{2}+A\bm{k}\cdot\bm{\sigma}+(M-Bk^{2})\sigma_{z},$ (1)
where $\bm{k}=(k_{x},k_{y})$ is the quasimomentum, and $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ are the Pauli matrices acting on the $E_{1}/H_{1}$ space of subbands of HgTe quantum wells [56]. As standard, we use $A=365$ meV nm, $B=686$ meV nm2, and the mass term $M=2$ meV, unless otherwise specified. For simplicity, we choose to set $C=D=0$ to keep the energy spectrum particle-hole symmetric, and to allow for a direct comparison with the Haldane [57] and Kane-Mele [58] models (see Supplemental Material [59] for details). For the numerical calculations, $\bm{k}=-i\bm{\nabla}$ is discretized into the square lattice with lattice parameter $a=3$ nm, defining finite systems of variable sizes $L_{x}\times L_{y}$. The Anderson disorder is introduced as uncorrelated random onsite energies uniformly distributed in the range $[-W/2,W/2]$. We compute the Landauer conductance $G$ and the current densities by standard methods [62, 63], considering this region coupled to semi-infinite pristine leads of width $L_{y}$.
Figures 1(a) and 1(b) show the band structure of the BHZ Hamiltonian in a ribbon geometry of width $L_{y}=600$ nm and periodic along $x$ for $M=+2$ meV and $M=-2$ meV, respectively. In both cases, we consider pristine systems ($W=0$) to emphasize that, in the trivial insulator case, there is a band gap of $\approx 2M=4$ meV, while in the nontrivial regime ($M<0$) the gap is filled with topologically protected helical edge states, which are responsible for a quantized conductance of $G_{0}=e^{2}/h$. As shown in Ref. [23], and revised in the Supplemental Materials [59], within the 1BA the Anderson disorder ($W\neq 0$) renormalizes the mass term as
$M\rightarrow\widetilde{M}\approx M-\alpha(M)W^{2}.$ (2)
This important result reveals that disorder can drive the system from a trivial regime ($M=2$ meV, $W=0$) into a TAI regime with an effective mass of $\widetilde{M}<0$. Figure 1(c) shows the conductance $G$ for the disordered BHZ model as a function of the Fermi energy $\varepsilon_{F}$ and the disorder intensity $W$. The large parameter space region of the quantized conductance $G=G_{0}$ (in yellow) characterizes a TAI phase. Figures 1(d) and 1(e) illustrate the current and spin densities for a smaller system ($L=100$ nm) for different disorder intensities and $\varepsilon_{F}=0$. For small $W$, as shown in Fig. 1(d), the system is already near a TAI phase and the current and spin densities show a tendency to form helical edge states. Figure 1(e) shows that as the disorder strength $W$ is increased, the system reaches the TAI phase and the helical edge states become well-defined. For even larger $W$ values, comparable with the bandwidth, the topological protection is destroyed and the states become localized, as shown by Fig. 1(f), indicating the onset of Anderson localization.
Local Chern marker.– As discussed in Ref. [22], in the trivial regime the conductance is dominated by bulk transport, and the scaled conductance $\sigma=GL_{x}/L_{y}$ decays with $W$ as a universal function, independent of the ribbon width $L_{y}$. In contrast, at the TAI phase, $G=G_{0}$ forms a quantized plateau [see Fig. 1(c)]. Therefore, instead of using $G$, we find it convenient to use local markers to investigate the normal-TAI transition, its scale invariance properties, and critical exponents. As shown in Ref. [29], the local Chern marker can be written as [64]
[…]
where $\hat{U}$ is defined below, the integral is taken over a unit cell of area $A_{c}$ centered at $\bm{R}$, the sum runs over the states $\beta$ that define the basis for the model Hamiltonian (e.g., $E_{1}/H_{1}$ orbitals for the BHZ model), $\hat{x}$ and $\hat{y}$ are the position operators, and $\hat{P}$ is the projector over occupied states, defined as
$\hat{P}=\sum_{\ell}^{\epsilon_{\ell}<\varepsilon_{F}}\outerproduct{\psi_{\ell}}{\psi_{\ell}}.$ | . (4)
Here $\epsilon_{\ell}$ and |$\ket{\psi_{\ell}}$ are the eigenenergies and eigenstates of the Hamiltonian $H$ for a finite, nonperiodic system, and the Fermi energy is set to $\varepsilon_{F}=0$ hereafter. It has been shown that $\mathcal{C}(\bm{R})$ recovers the quantized bulk Chern number $C$ as one restores the lattice periodicity for clean systems [31], while for finite systems $\mathcal{C}(\bm{R})\approx C$ for $\bm{R}\approx 0$, near the center of the sample [29, 21]. For the BHZ model, defined by Eq. (1), and for the Haldane model [59], we consider $\hat{U}=1$. On the other hand, for the spinful Kane-Mele model [59], we set $\hat{U}=\sigma_{z}$ (Pauli matrix in spin space), such that $\mathcal{C}(\bm{R})$ represents a local spin-Chern number [65]. In this case, for a finite Rashba coupling, discussed in Ref. [59], the spin-Chern number is not quantized, but it remains a reliable witness for nontrivial topology, as shown in Refs. [66, 67, 68].
In the presence of Anderson disorder, the local marker $\mathcal{C}(\bm{R})$ acquires fluctuations. Consequently, the analysis of the LCM requires a statistical approach. Here, we focus on the average $\mathcal{C}(\bm{R})$, which can be taken over disorder realizations and/or spatial averages. For large systems, this can be an intensive numerical task. Therefore, we have implemented two approaches. First, to calculate full $\mathcal{C}(\bm{R})$ maps as a function of $\bm{R}$ for a single-disorder sample (e.g., in Fig. 2), we use a brute force approach, where the finite-size system is fully diagonalized, $H\ket{\psi_{\ell}}=\epsilon_{\ell}\ket{\psi_{\ell}}$, such that the operators $\hat{P}$, $\hat{x}$, and $\hat{y}$ can be explicitly calculated to yield $\mathcal{C}(\bm{R})$ from Eq. (3). In the second approach, we define the disorder-averaged local marker $\mathcal{C}_{0}$ as the average of $\mathcal{C}(\bm{R}=0)$ over many disorder realizations, namely, $\mathcal{C}_{0}\equiv\langle\mathcal{C}(\bm{R}=0)\rangle$. We compute $\mathcal{C}_{0}$ employing the efficient implementation developed in Ref. [69] based on the kernel polynomial method (KPM) [70]. The KPM allows us to estimate $\hat{P}\ket{\varphi}$ for an arbitrary state |$\ket{\varphi}$ in terms of an efficient expansion of Chebyshev polynomials (details about the KPM implementation are presented in Ref. [59]).
Figure 2 shows $\mathcal{C}(\bm{R})$ as a function of $\bm{R}=(X,Y)$ for a single-disorder realization for increasing disorder strengths $W$. Figure 2(a) corresponds to a small $W$ for which the system is in the trivial regime, which is characterized by a small LCM near the sample center, that is, $\mathcal{C}({\bm{R}}\approx 0)\approx 0$. Figure 2(b) considers a moderate disorder intensity $W$, yielding $\mathcal{C}({\bm{R}}\approx 0)\approx 0.7$ and thus approaching a TAI regime, which is fully achieved for larger $W$, as shown in Fig. 2(c), where $\mathcal{C}({\bm{R}}\approx 0)\approx 1.0$. Notice that Figs. 2(b) and 2(c) correspond to the same values of $W$ of Figs. 1(d) and 1(e). Interestingly, up to Fig. 2(c) or Fig. 1(e), we find that the Anderson disorder has driven the system into a TAI phase, but there are no clear signs of Anderson localization, since the maps are overall smooth, apart from small fluctuations that cause some blurriness. However, for even larger $W$ [see Fig. 2(d) or Fig. 1(f)], Anderson localization emerges as strong fluctuations over the full map $\mathcal{C}(\bm{R})$. In all cases, we see that $\mathcal{C}(\bm{R})$ takes negative values at the edges, which fulfills the constraint $\sum_{\bm{R}}\mathcal{C}(\bm{R})=0$ [29]. Indeed, this is an undesirable consequence of the approximations that take place in the derivation of the local marker [64], which, for now, we leave as an issue to be discussed in future works.
Figure 2: Local Chern marker $\mathcal{C}(\bm{R})$ for a representative single-disorder realization as a function of $\bm{R}=(X,Y)$ for a system of size $L_{x}=L_{y}=100$ nm, $M=+2$ meV, and disorder strengths (a) $W=25$ meV, (b) $W=180$ meV, (c) $W=400$ meV, and (d) $W=897$ meV. Panels (b) to (d) correspond to the current and spin densities shown in Figs. 1(e) and 1(f).
Phase transition and scaling.– Figure 3 shows the disorder-averaged local Chern marker $\mathcal{C}_{0}$ as a function of $W$ for different system sizes in a square geometry, $L_{x}=L_{y}\equiv L$, and a fixed $M=+2$ meV. As $W$ increases, all cases first display a transition from a trivial regime with $\mathcal{C}_{0}\approx 0.1$ to a nontrivial regime characterized by $\mathcal{C}_{0}\approx 1$. For very large values of $W$, exceeding the bandwidth ($E_{\rm max}\sim|B|k_{\rm max}^{2}=8|B|/a^{2}\approx 610$ meV for $a=3$ nm) of the pristine BHZ Hamiltonian, $\mathcal{C}_{0}$ collapses back to $\mathcal{C}_{0}\rightarrow 0$. This is the transition seen from Fig. 2(c) to Fig. 2(d), which is due to the emergence of Anderson localization. Such behavior is not surprising, since in such a strong-disorder regime, it hardly makes sense to consider the system as topological. Figure 3(a) clearly shows that the second transition, $\mathcal{C}_{0}=1\rightarrow 0$, does not depend on $L$. In fact, the transition to a trivial Anderson insulator is governed by an energy scale defined by the bandwidth [71], which is only influenced by $L$ for very small samples. Thus, for moderate $L$, both the bandwidth and this transition are expected to be independent of $L$. In contrast, the first transition, $\mathcal{C}_{0}\approx 0.1\rightarrow 1$, is $L$ dependent. This dependence is what we analyze next.
We obtain a single-parameter universal scaling function for the trivial-TAI phase transition by expressing ${\cal C}_{0}$ in terms of the parameter $z$, which combines the disorder strength $W$ and the system size $L$ according to
$\mathcal{C}_{0}(W,L)\rightarrow\mathcal{C}_{0}(z),$ (5)
with
$z=\dfrac{W^{\mu}-W_{c}^{\mu}}{W_{\rm max}^{\mu}}\cdot\dfrac{L}{L_{\rm max}},$ (6)
as shown in Fig. 3(b). Here, $W_{c}$ is the critical point where the curves corresponding to systems with different sizes $L$ cross, and $\mu$ is the critical exponent. To make $z$ dimensionless, we introduce $L_{\rm max}=250$ nm, the largest simulated system size, and $W_{\rm max}=1200$ meV, the largest considered disorder strength, in the denominators of $z$. We find numerically that the optimal scaling in Fig. 3(b) occurs for $\mu\approx 2$ and $W_{c}\approx 86$ meV [dashed line in Fig. 3(a)]. Next, to understand these values for $\mu$ and $W_{c}$, we analyze the phase diagram and scaling properties of the $\mathcal{C}_{0}$ as a function of both the mass $M$ and the disorder intensity $W$.
Figure 3: (a) Disorder-averaged local Chern marker (400 disorder realizations and 2000 Chebyshev moments in the KPM expansion) as a function of $W$ for different $L_{x}=L_{y}=L$ showing a trivial-TAI transition, $\mathcal{C}_{0}\rightarrow 0$ to 1, followed by the trivial Anderson localization as $W$ becomes larger than the bandwidth. (b) Optimal scaling of the marker showing that all curves in panel (a) fall into a universal trend for the trivial-TAI transition region upon the scaling $W\rightarrow z=(W^{\mu}-W_{c}^{\mu})L/W_{\rm max}^{\mu}L_{\rm max}$, with $\mu\approx 2$ and $W_{c}\approx 86$ meV. The points indicated by the arrows (i) to (iv) correspond to the ${\cal C}({\bm{R}})$ color maps shown in Fig. 2(a) to 2(d), respectively.
Phase diagram.– The local Chern marker can be used to define a topological phase diagram. For the BHZ model, we now consider $\mathcal{C}_{0}$ in terms of the mass $M$ and the disorder strength $W$, as shown in Fig. 4, for two system sizes. In both cases the color code labels the value of the local marker $\mathcal{C}_{0}$ calculated for 400 disorder realizations. The phase diagrams in Figs. 4(a) and 4(b) show the following phases: trivial, TI, TAI, and Anderson localization (AL). As discussed above, the AL phase occurs when the disorder strength exceeds the bandwidth $W>E_{\rm max}\approx 610$ meV.
We recall that, for $W=0$, the bulk Chern number indicates that a TI phase occurs for $M<0$, whereas Figs. 4(a) and 4(b) show that the trivial to TI transition, characterized by the local marker $\mathcal{C}_{0}$, depends upon the finite system size $L$. In a pristine 2D TI, the topological edge states’ are gapless and do not hybridize, provided the system size $L$ is much larger than the edge states penetration depth $\xi=A/|M|$, i.e., for $A/|M|\ll L$ (with $M<0$ and $W=0$). We extend this gapless condition to the disordered case by making $M\rightarrow\widetilde{M}$ to obtain an expression for the phase boundary using $\widetilde{M}$ from 1BA in Eq. (2). Hence, from the gapless condition $A/|\widetilde{M}|\ll L$ and assuming $\widetilde{M}\leq 0$, we obtain
$W_{\rm gapless}(M)=\sqrt{\dfrac{M+A/L}{\alpha(M)}}\approx\sqrt{\dfrac{M+A/L}{\alpha_{0}}}.$ (7)
Here and in what follows we use that the first Born approximation gives $\alpha(M)\approx\alpha_{0}$ [59]. In Figs. 4(a) and 4(b) the dashed lines mark the phase boundary given by $W_{\rm gapless}(M)$ for an optimal $\alpha_{0}\approx 1/3700$ meV-1, which matches remarkably well the phase boundary between the trivial and topological phases (TI and TAI).
Figure 4: (a, b) Phase diagrams calculated using the local Chern marker $\mathcal{C}_{0}$ as a function of the BHZ mass $M$ and the disorder strength $W$ for (a) $L=50$ nm and (b) 100 nm. In all cases, we consider 200 disorder realizations and 1000 Chebyshev moments on the KPM expansion. The dashed white line marks the phase boundary between the trivial and topological phases (TI and TAI) obtained from the gapless condition $A/|\widetilde{M}|\ll L$. The dot-dashed line marks the boundary where $W$ matches the bandwidth of the clean system, above which the Anderson localization phase takes place. (c) All data from panels (a) and (b), and additional data for a system of $L=150$ nm, fall into a universal curve for the trivial to TAI phase transition upon the scaling $(M,W,L)\rightarrow z=[W^{\mu}-W_{c}^{\mu}(M)]L/W^{\mu}_{\rm max}L_{\rm max}$.
The data shown in Figs. 4(a) and 4(b) for all $M$ and $W$, and additional data for $L=150$ nm (not shown), can be combined into a single universal function that characterizes the trivial to TAI phase transition shown in Fig. 4(c). The scaling used in this case is a direct extension of the one presented in Eqs. (5) and (6), where now $W_{c}\rightarrow W_{c}(M)$ depends upon $M$. To obtain an expression for $W_{c}(M)$ and to understand the value $\mu\approx 2$ of the critical parameters, let us first consider the clean regime ($W=0$), to find the scaling law for the trivial to TI phase transition as a function of $M$. Here, we consider an argument similar to the one used in Ref. [21] for the Haldane model, as follows. In the topological regime, the only relevant length scales are $\xi=A/|M|$ and $L$, as introduced above. Therefore, one can expect that the topological phase transition scaling with the system size $L$ should be a universal function of $L/\xi\propto ML$ (assuming $A$ is constant). Thus, we justify that
$\mathcal{C}_{0}(M,L)\equiv\mathcal{C}_{0}(ML).$ (8)
Next, to extend this scaling to finite $W$, we replace $M\rightarrow\widetilde{M}$ using the 1BA expression from Eq. (2) with $\alpha(M)\approx\alpha_{0}$, such that the universal function argument now reads
$\widetilde{M}L=(M-\alpha_{0}W^{2})L=-\alpha_{0}(W^{2}-W_{c}^{2})L,$ (9)
where $W_{c}^{2}=M/\alpha_{0}$ is the critical point and marks the phase boundary in the thermodynamic limit ($L/\xi\rightarrow\infty$), where $\widetilde{M}=0$ and $W_{\rm gapless}=W_{c}$. Indeed, for $M=2$ meV, this expression yields $W_{c}\approx 86$ meV, thus matching the value found in Fig. 3. Moreover, apart from constant factors, the expression for $\widetilde{M}L$ in Eq. (9) matches $z$ given by Eq. (6) with $\mu=2$, in agreement with the previously numerically obtained value.
Conclusions.– We have investigated the scale invariance of the local Chern marker $\mathcal{C}_{0}$ applied to TAIs. The smooth profile of $\mathcal{C}_{0}$ throughout the phase transition allows us to characterize the scale invariance and obtain critical parameters, and universal scaling functions. Our analysis uses the disordered BHZ model as a canonical example of TI and TAI phases. The simplicity of the model allows for an analytical justification of the obtained numerical results. We like to stress that the disordered Haldane [57] and Kane-Mele [58] models exhibit an identical scaling and universal behavior, as we have shown in the Supplemental Material [59]. This observation supports the statement that our findings are quite generic. Indeed, as long as the first-order Born approximation remains sufficiently accurate, we expect that the universal function will scale with $\sim W^{2}L$, while deviations from this behavior should occur only beyond the validity range of the 1BA. We believe that our findings shed light on an approach to investigating scaling properties of topological phase transitions for disordered and amorphous systems.
Note added in proof. The independent study [72] has considerable overlap with our work.
Acknowledgments.– This work was supported by the Brazilian funding agencies CNPq, CAPES, FAPERJ, and FAPEMIG (Grant No. PPM-00798-18). The authors acknowledge useful conversations with C. Beenakker, A. Akhmerov, and C. P. Orth.